# Reconstruction of Rough Conductivities from Boundary Measurements

### v3

@@ 1. Introduction
An interesting problem that has received considerable interest is of finding the minimum regularity assumptions on $\gamma$ (or $q$) under which injectivity and the reconstruction procedure hold. This question is also of practical importance. For example, it was pointed out in [18] that if $q$ arises from a Gaussian random field satisfying certain conditions, almost every instantiation of $q$ belongs to a Sobolev space of fixed negative order. For $n\geq 3$, the regularity assumption for uniqueness was relaxed to $\gamma\in C^{3/2+}$ in [9], to $C^{3/2}$ in [44], to $W^{3/2,2n+}$ in [10], to $W^{3/2+,2}$ in [43] and to $\gamma\in C^{1}$ or $\gamma\in C^{0,1}$ with $\|\nabla\log\gamma\|_{L^{\infty}}$ small in [24]. The smallness condition was removed in [19]. Note that for $\gamma\equiv 1$ in a neighborhood of $\partial\Omega$ (as we assume in this paper), the regularity assumption in [10] is reduced to $\gamma\in W^{3/2,2n}$. It was also conjectured by Brown in [10] that uniqueness holds for $\gamma\in W^{1,n}$ for all $n\geq 3$. This was proved for $n=3,4$ in [23].
For the problem of reconstruction, Nachman’s procedure in [41] was adapted to the case of $\gamma\in C^{1}$ or $\gamma\in C^{0,1}$ with $\|\nabla\log\gamma\|_{L^{\infty}}$ sufficiently small in [21]. The approaches in [43] and [23] present technical difficulties for reconstruction since the bounds on the multiplication operator are only established for large $|\zeta|$ on average, and the approach in [19] produces only local solutions. In this paper, we prove the validity of Nachman’s reconstruction procedure for $\gamma\in W^{3/2,2n}$ with $\gamma\equiv 1$ near $\partial\Omega$ as in [10]. The key step is establishing bounds on the multiplication operator $\varphi\mapsto q\varphi$ in the weighted Sobolev spaces of Sylvester and Uhlmann [46], similar to the bounds in [18]. Note that functions in $W^{3/2,2n}$ need not be Lipschitz. However, $W^{3/2,2n}$ is contained in the Zygmund space $C^{1}_{*}$ of continuous functions $f$ such that
@@
We refer the reader to the monograph [47] for more on Zygmund spaces. We also note here that since the original version of this paper was first posted on arXiv, reconstruction results have also been proved for Lipschitz conductivities by Caro, García-Ferrero, and Rogers [17].
Another question of interest is of stability of the map $\gamma\mapsto\Lambda_{\gamma}$. It was shown by Alessandrini in [1] that under the a-priori assumption
@@
where $\sigma=\sigma(n,s)\in(0,1)$. Subsequently, the a-priori assumptions were relaxed to $\|\gamma_{j}\|_{W^{2,\infty}}\leq M$ in [2, 3]. Such logarithmic estimates were shown to be optimal up to the value of the exponent by Mandache in [39] via explicit examples. Later, stability was proved for conductivities bounded a-priori in $C^{1,\frac{1}{2}+\epsilon}\cap H^{n/2+\epsilon}$ with $\partial\Omega$ smooth by Heck in [25] and for a-priori bounds in $C^{1,\epsilon}(\overline{\Omega})$ with $\partial\Omega$ Lipschitz by Caro, García and Reyes in [16]. In this paper, we prove a similar log-type stability estimate with $\|\gamma_{j}\|_{W^{2-s,n/s}(\Omega)}\leq M$ for some $0<s<1/2$. The main results of this paper are summarized in the following theorem:
@@ 5. Stability
Let $0<s<1/2$ and $q_{1},q_{2}\in W^{-s,n/s}_{\textrm{comp}}(\Omega)$ satisfy the a-priori estimate
@@
As in Proposition 2.3, let us extend $\gamma_{j}$ to all of $\mathbb{R}^{n}$ by defining $\gamma_{j}\equiv 1$ on $\mathbb{R}^{n}\setminus\Omega$, so that $\gamma_{j}-1\in W^{-s,n/s}_{\textrm{comp}}(\Omega)$. Note that this implies $\gamma_{j}\in C^{1+\epsilon}_{*}=C^{1,\epsilon}$ for $\epsilon=1-2s$. Define $q_{j}=\gamma_{j}^{-1/2}\Delta\gamma_{j}^{1/2}$. Also choose a bounded domain $U$ such that $\overline{\Omega}\subset U$ and $\partial U$ is smooth. We observe that the function $v=\log\gamma_{1}-\log\gamma_{2}$ solves the following elliptic boundary value problem:
@@ Acknowledgments
The author would like to thank Gunther Uhlmann and Hart Smith for their helpful suggestions and comments on the manuscript, as well as their mentorship throughout this project. The project was completed during the author’s tenure as a Graduate Teaching Assistant in the Department of Mathematics at the University of Washington.

### v4

@@ 1. Introduction
An interesting problem that has received considerable interest is of finding the minimum regularity assumptions on $\gamma$ (or $q$) under which injectivity and the reconstruction procedure hold. This question is also of practical importance. For example, it was pointed out in [18] that if $q$ arises from a Gaussian random field satisfying certain conditions, almost every instantiation of $q$ belongs to a Sobolev space of fixed negative order. For $n\geq 3$, the regularity assumption for uniqueness was relaxed to $\gamma\in C^{3/2+}$ in [9], to $C^{3/2}$ in [44], to $W^{3/2,2n+}$ in [10], to $W^{3/2+,2}$ in [43] and to $\gamma\in C^{1}$ or $\gamma\in C^{0,1}$ with $\|\nabla\log\gamma\|_{L^{\infty}}$ small in [24]. The smallness condition was removed in [19]. It was also conjectured by Brown in [10] that uniqueness holds for $\gamma\in W^{1,n}$ for all $n\geq 3$. This was proved for $n=3,4$ in [23].
For the problem of reconstruction, Nachman’s procedure in [41] was adapted to the case of $\gamma\in C^{1}$ or $\gamma\in C^{0,1}$ with $\|\nabla\log\gamma\|_{L^{\infty}}$ sufficiently small in [21]. In this paper, we extend Nachman’s reconstruction procedure to the Sobolev scale $W^{3/2,2n}$ for conductivities that are identically $1$ near $\partial\Omega$. Note that functions in $W^{3/2,2n}$ need not be $C^{1}$ or even Lipschitz, but do belong to the Zygmund space $C^{1}_{*}$ (c.f. [47]) of continuous functions $f$ such that
@@
To the best of our knowledge, this paper presents the first constructive result for the Calderón problem in dimensions $n\geq 3$ for a class of conductivities that includes non-Lipschitz functions. We also note here that since the original version of this paper was first posted on arXiv, reconstruction results have also been proved for Lipschitz conductivities by Caro, García-Ferrero, and Rogers [17].
Our proof avoids the technical obstruction noted in [10] regarding Sobolev extensions of $\gamma$ to $\mathbb{R}^{n}$ by working only with conductivities that are equal to $1$ near the boundary. While the reduction to the Schrodinger equation for $\gamma\in W^{3/2,2n}$ produces $q\in W^{-1/2,2n}$ (which was treated in [10] in the context of uniqueness), the validity of this reduction for the purposes of reconstruction, in the sense of equality of the corresponding Dirichlet-to-Neumann maps (see Proposition 2.3) had not been established before at this level of regularity.
Another question of interest is of stability of the map $\gamma\mapsto\Lambda_{\gamma}$. It was shown by Alessandrini in [1] that under the a priori assumption
@@
where $\sigma=\sigma(n,s)\in(0,1)$. Subsequently, the a priori assumptions were relaxed to $\|\gamma_{j}\|_{W^{2,\infty}}\leq M$ in [2, 3]. Such logarithmic estimates were shown to be optimal up to the value of the exponent by Mandache in [39] via explicit examples. Later, stability was proved for conductivities bounded a priori in $C^{1,\frac{1}{2}+\epsilon}\cap H^{n/2+\epsilon}$ with $\partial\Omega$ smooth by Heck in [25] and for a priori bounds in $C^{1,\epsilon}(\overline{\Omega})$ with $\partial\Omega$ Lipschitz by Caro, García and Reyes in [16]. In this paper, we prove a similar log-type stability estimate with $\|\gamma_{j}\|_{W^{2-s,n/s}(\Omega)}\leq M$ for some $0<s<1/2$. In the Hölder scale, this corresponds to $C^{1,\epsilon}$ as in [16], with $\epsilon=1-2s$. However, our use of the Sobolev scale allows us to avoid Bourgain-type spaces, leading to a much simpler and shorter proof.
The main results of this paper are summarized in the following theorem:
@@ 5. Stability
Let $0<s<1/2$ and $q_{1},q_{2}\in W^{-s,n/s}_{\textrm{comp}}(\Omega)$ satisfy the a priori estimate
@@
As in Proposition 2.3, let us extend $\gamma_{j}$ to all of $\mathbb{R}^{n}$ by defining $\gamma_{j}\equiv 1$ on $\mathbb{R}^{n}\setminus\Omega$, so that $\gamma_{j}-1\in W^{2-s,n/s}_{\textrm{comp}}(\Omega)$. Note that this implies $\gamma_{j}\in C^{1+\epsilon}_{*}=C^{1,\epsilon}$ for $\epsilon=1-2s$. Define $q_{j}=\gamma_{j}^{-1/2}\Delta\gamma_{j}^{1/2}$. Also choose a bounded domain $U$ such that $\overline{\Omega}\subset U$ and $\partial U$ is smooth. We observe that the function $v=\log\gamma_{1}-\log\gamma_{2}$ solves the following elliptic boundary value problem:
@@ Acknowledgments
The author would like to thank Gunther Uhlmann and Hart Smith for their helpful suggestions and comments on the manuscript, as well as their mentorship throughout this project. This project was initiated and largely completed during the author’s tenure as a Graduate Teaching Assistant in the Department of Mathematics at the University of Washington.